OpenStudy (anonymous):

It takes a lion 4 hours to eat a sheep. A leopard needs 5 hours and a wild dog needs 6 hours to eat a sheep. How long will it take the three of them together to eat a Sheep? Please Help

OpenStudy (phi):

Do you need a recipe for lamb chops?

OpenStudy (anonymous):

I need the answer Please Phi

OpenStudy (anonymous):

1?4+1/5+1/6=1/x

OpenStudy (anonymous):

x=60/37=1.62 hours

OpenStudy (anonymous):

Solve the following for hours x by considering how much sheep each can animal eat in an hour: \[\left(\dfrac{1}{4} \cdot \dfrac{\text{sheep}}{\text{hour}}\right) + \left(\dfrac{1}{5} \cdot \dfrac{\text{sheep}}{\text{hour}}\right) + \left(\dfrac{1}{6} \cdot \dfrac{\text{sheep}}{\text{hour}}\right)=\left(\dfrac{1}{x} \cdot \dfrac{\text{sheep}}{\text{hour}}\right)\]

OpenStudy (anonymous):

I need the answer Please Phi

OpenStudy (anonymous):

Yakeyglee so what is the answer?

OpenStudy (phi):

I think you have to add 1/4 plus 1/5 plus 1/6 and then flip the answer